.a. Define elasticity of supply and find the price from the given statement:

If Es of a good is 2 and a firm supplies 200 units at price of Rs 8 per unit, then at what price will the firm supply 250 units.

.b. Calculate the elasticity of supply if a 15 %increase in the price of soya bean oil increases its supply from 300 to 345 units


1
Expert's answer
2021-05-13T11:01:54-0400
"E_S=\frac{(Q_2-Q_1)(p_2+p_1)}{(Q_2+Q_1)(p_2-p_1)}"

"p_2=8.9"

"E_d=\frac{\frac{345}{300}}{1.15}=1"